OpenStudy (anonymous):

supplementary angles with measures and degrees 6x-4 and 8x-12 degrees

OpenStudy (katrinakaif):

If they are supplementary, then they have to be equal to..?

OpenStudy (anonymous):

180

OpenStudy (katrinakaif):

Since these two measurements are supplementary, set both of them equal to 180

OpenStudy (katrinakaif):

Therefore, (6x - 4) + (8x - 12) = 180

OpenStudy (anonymous):

so I can just add them together and use algebra to find x? 14x-16=180 14x=164 x = 11.7142857142857

OpenStudy (katrinakaif):

Now, x is equal to 11.71. You have to plug it back into the given measurements as "x" to get their degrees

OpenStudy (anonymous):

sorry did it wrong lol

OpenStudy (katrinakaif):

No no, you're on the right track. Just needed to fulfill some additional steps. Great work! ;)

OpenStudy (anonymous):

it's 14x-16=180 14x = 196 x = 14

OpenStudy (katrinakaif):

Ah there we go. My mistake I did not point that out! =P Now, you can plug them into the original measurements to conclude to your final answer ;)

OpenStudy (anonymous):